OpenStudy (anonymous):

Divide and simplify. 3w-12/10 divided by w-4/14w

OpenStudy (anonymous):

Thanks!

OpenStudy (campbell_st):

\[\frac{3w -12}{10} \times \frac{14w}{w -4} ... =\frac{3(w - 4) \times 14w}{10(w -4)}... = 21w/5\]

OpenStudy (anonymous):